OpenStudy (thomasfire55):

The sum of the roots of 3x^2 +11x-4=0 is

OpenStudy (thomasfire55):

a) -11/3 b) 11/3 c) -4/3 d) 4/3

ganeshie8 (ganeshie8):

Suppose the roots are a and b

ganeshie8 (ganeshie8):

Then we can factor the quadratic as 3x^2 +11x-4 = 3(x-a)(x-b)

ganeshie8 (ganeshie8):

Same as 3x^2 +11x-4 = 3[x^2 - (a+b)x +ab]

OpenStudy (thomasfire55):

?

ganeshie8 (ganeshie8):

comparing x coefficients both sides gives 11 = -3(a+b)

ganeshie8 (ganeshie8):

You must know what to do next to get the sum of roots a+b

OpenStudy (thomasfire55):

no i never paid attetion in math

ganeshie8 (ganeshie8):

Its never too late for learning anything

OpenStudy (thomasfire55):

can you teach me what to do

ganeshie8 (ganeshie8):

It will be easy if you take time to understand the question I was asking

OpenStudy (thomasfire55):

still confussed

ganeshie8 (ganeshie8):

11 = -3(a+b) Can you tell me what I was asking you to do after this step ?

ganeshie8 (ganeshie8):

Your goal is a+b, so simply divide -3 both sides and get -11/3 = a+b

OpenStudy (thomasfire55):

so its -11/3?

ganeshie8 (ganeshie8):

Yes. If I may give a piece of advice. It can be confusing, but it will make sense if you persist and fi a few more practice problems. Don't simply give up just because youve felt it confusing the first time!